OpenStudy (anonymous):

A rock is launched from a cannon. After 1 second, the rock is 243 feet in the air; after 2 seconds, it is 452 feet in the air. Using the quadratic function, find the height, in feet, of the rock after 5 seconds in the air

OpenStudy (anonymous):

I've seen a similar problem to it yesterday, and I got confused lol ^_^"

OpenStudy (anonymous):

What is the angle at which it is launched?

OpenStudy (anonymous):

doesn't say.

OpenStudy (amistre64):

a+b = 243 a = 243 - b 4a +2b = 452 4(243-b) +2b = 452 892 -4b + 2b = 452 892 -452 = 2b 440/2 = b b = 220 -------------- a = 243-220 = 23 23x^2 +220x = y -------------------- 23 + 220 = 243 checks 23(4) + 220(2) = 452 92 + 440 = 30 ok.....whatd I do wrog..hard to type in the dark :)

OpenStudy (anonymous):

assume it is launched straight up. say its initial acceleration is a. distance is ut+1/2at^2

OpenStudy (anonymous):

initial speed is 0. so d = 1/2 at^2

OpenStudy (amistre64):

its a simple quad equation, not physics....but that might be a shot :)

OpenStudy (amistre64):

(0,0) (1,243) (2,452)

OpenStudy (anonymous):

since it is launched straight up in the air, the gravity is acting against the rock. so s = 1/2(a-32.17)t^2

OpenStudy (amistre64):

1a + 1b = 243 b = 243-a 4a + 2b = 452 4a + 2(243-a) = 452 4a +486 -2a = 452 2a = 452 - 486 2a = -34 a = -17 by that

OpenStudy (anonymous):

oh okay, then you should follow amistre 64's procedure.

OpenStudy (amistre64):

b = 243 --17 b = 260 -17x^2 +260x = y -17(4) + 260(2) = 452 -68 + 520 = 452 right? someone tell me its right lol

OpenStudy (anonymous):

452 is the height of the rock at 2 seconds.

OpenStudy (anonymous):

amistre64 can u please help me

OpenStudy (amistre64):

the numbers are working yay!!...plug in a 5

OpenStudy (anonymous):

plug in a 5 where?

OpenStudy (amistre64):

-17(25) +260(5) = ?

OpenStudy (anonymous):

that's what i thought. i just tried it. i got 875

OpenStudy (amistre64):

-425 + 1300= 875....me too :)

OpenStudy (anonymous):

correct! one more to go! let's hope i pass it this time! i've only missed one outta 4. on to the 5th question. thanks amistre64!

OpenStudy (amistre64):

i wrote you a quad program today, but left my flash drive at the college lol

OpenStudy (anonymous):

thanks though!